# Privacy Preserving Federated Unsupervised Domain Adaptation with Application to Age Prediction from DNA Methylation Data

### v1

@@ Abstract
In computational biology, predictive models are widely used to address complex tasks, but their performance can suffer greatly when applied to data from different distributions. The current state-of-the-art domain adaptation method for high-dimensional data aims to mitigate these issues by aligning the input dependencies between training and test data. However, this approach requires centralized access to both source and target domain data, raising concerns about data privacy, especially when the data comes from multiple sources. In this paper, we introduce a privacy-preserving federated framework for unsupervised domain adaptation in high-dimensional settings. Our method employs federated training of Gaussian processes and weighted elastic nets to effectively address the problem of distribution shift between domains, while utilizing secure aggregation and randomized encoding to protect the local data of participating data owners. We evaluate our framework on the task of age prediction using DNA methylation data from multiple tissues, demonstrating that our approach performs comparably to existing centralized methods while maintaining data privacy, even in distributed environments where data is spread across multiple institutions. Our framework is the first privacy-preserving solution for high-dimensional domain adaptation in federated environments, offering a promising tool for fields like computational biology and medicine, where protecting sensitive data is essential.
@@
Federated Learning, Unsupervised Domain Adaptation, Data Privacy
@@ 1 Introduction
Machine learning (ML) has rapidly become a powerful tool with applications across numerous fields, including computational biology and healthcare, where it has demonstrated significant potential in solving complex problems [1, 2, 3, 4]. Motivated by the increasing availability of data and advancements in computational hardware, predictive models are now capable of discovering new relationships, such as those between genotypes and phenotypes [5], and improving healthcare [6]. Today, these systems assist healthcare professionals in predicting critical events, diagnosing diseases, and forecasting patient responses to treatments, thereby enhancing patient outcomes.
ML models are traditionally built on the assumption that the distribution of the training data matches that of the data the model will later encounter. This assumption forms the basis for most ML applications. However, in real-world scenarios, this usually cannot be guaranteed. For instance, different hospitals might use different experimental equipment, protocols and might treat very different patient populations, which can lead to distributional differences between the training and the test data [7]. Moreover, if new data deviates from the training distribution, the learned relationships may no longer be valid, causing significant declines in model performance [6].
Domain adaptation is a specialized area in ML, designed to develop predictive models that perform well even with distribution mismatches between training and new data [8]. In domain adaptation, the training and test sets are referred to as the source and target domains, respectively. The main goal is to create a model that generalizes effectively to the target domain while being primarily trained on data from the source domain. Various domain adaptation methods exist [9], each differing in how much information from the target domain is used to reduce distribution differences and improve prediction accuracy. Supervised domain adaptation techniques, which require labeled samples from the target domain, address domain mismatch by emphasizing target domain samples during training. In contrast, unsupervised domain adaptation, a more challenging variant, only has unlabeled data from the target domain available for training. This scenario lacks trivial methods for including target domain samples during training or assessing the model’s performance post-training. Unsupervised domain adaptation approaches, especially those based on deep neural networks, have shown high performance in computer vision [10, 11]. Despite the success of deep learning, its application in computational biology often requires alternative approaches due to the need for model explainability, typically limited data, and high dimensionality. In computational biology, the number of features often greatly exceeds the number of samples, making deep learning less feasible. Furthermore, most of the existing unsupervised domain adaptation techniques focus on classification problems [10, 11, 12], leaving unsupervised domain adaptation for regression tasks less explored and understudied.
Currently, the state-of-the-art solution for performing unsupervised domain adaptation on high-dimensional tabular datasets is a method called wenda (weighted elastic net for domain adaptation), which is based on a weighted elastic net [6]. Wenda addresses these issues by examining the dependency structure between inputs in both source and target domain data, penalizing features that show discrepancies while encouraging the model to focus on robust features with consistent behavior across domains. In the original paper, wenda was applied to the problem of age prediction from DNA methylation data across various human tissues [6]. While wenda significantly enhances performance compared to non-adaptive models, it has a notable limitation regarding data privacy, as it assumes that both target and source domain data are accessible simultaneously. This assumption renders it unsuitable for many real-world scenarios on human data where data is distributed across multiple parties that cannot openly share sensitive information.
As machine learning becomes increasingly popular, concerns about data privacy and confidentiality have led to increased regulatory focus, such as through the GDPR [^1] General Data Protection Regulation, European Union and CCPA [^2] California Consumer Privacy Act. This shift has increased interest in privacy-preserving techniques like Federated Learning (FL). FL involves multiple parties collaboratively training a machine learning model through a central server, such as a service provider or aggregator, while keeping their local data private. Although FL was initially proposed for cross-device applications, its relevance has grown to include other domains, such as collaborations between organizations or hospitals, where fewer but more reliable participants are involved.
In this context, we propose a new framework called freda (federated domain adaptation). By utilizing randomized encoding [13], secure aggregation [14], and federated learning[15], freda enables data owners in a distributed setting to perform privacy-preserving federated unsupervised domain adaptation on high-dimensional tabular data. Freda operates on the same underlying principles as wenda but removes the critical assumption of wenda, which requires simultaneous access to both the source and target domains. We evaluate our method on real-world data, focusing on the same problem of age prediction using DNA methylation data across multiple tissues. The key metric for assessing the success of domain adaptation on this data set is the model’s performance on the cerebellum samples, which are not represented in the training data and are known to be biologically distinct, even from other brain tissues, in terms of function and gene expression patterns [16, 17]. Our empirical results demonstrate that our framework can achieve a comparable level of performance to wenda on cerebellum samples, even when the training data is distributed across multiple sources, while maintaining complete data privacy. For the other tissues a baseline model without domain adaptation performs reasonably well. Our results show that the performance of our method remains consistent with up to 8 source clients, demonstrating that privacy-preserving unsupervised domain adaptation on high-dimensional omics data is possible even when training data is distributed among multiple entities.
@@ 2 Results
We use real data to showcase the performance of our proposed framework and how it compares to our baselines, specifically we focus on the problem of age prediction from DNA methylation data across multiple tissues.
@@ 2.1 Overview
We assume a scenario where the source domain is distributed across multiple input parties, and one input party holds unlabeled samples from the target domain. The goal is to perform unsupervised domain adaptation on the target domain using labeled data from the source domains, while preserving the data privacy of each contributing party.
Figure 1: Overview of the freda framework, in which multiple source domain clients collaborate to perform domain adaptation on the unlabeled target domain data with the assistance of an aggregator.
Freda operates with the same underlying principles as wenda. The main difference is that wenda has a central data repository, while freda operates on distributed data. Initially, the input dependency structure is collaboratively learned in a distributed setting by training feature models. This is followed by the computation of a confidence score for each feature and, finally, the collaborative training of the adaptive model on the source domains.
A high-level overview of freda is shown in Figure 1, illustrating the main phases of our framework. The protocol begins with training the feature models, which involves federated hyper-parameter optimization (1) followed by the computation of the closed-form solutions of the feature models (2). In the feature model training phase each feature in the dataset is modeled separately. For each feature, all clients split their data into two parts, removing one feature at a time to predict the excluded feature based on the remaining features. This step involves training multiple Gaussian process regressors, and is completed once all closed-form solutions for the feature models are computed using secure aggregation and the FLAKE randomized encoding framework [13] (Section 5.1).
Next, the aggregator sends the predictions of the feature models to the owner of the target domain for confidence score computation (3). The target client transforms these confidence scores into feature weights for each tissue and sends the feature weights to the aggregator, which then distributes them to the source clients.
Subsequently, the source clients, with the help of the aggregator, collaboratively train multiple federated weighted elastic nets with varying regularization parameters $\lambda$ (4). The trained models are sent to the target client, where the target client performs optimal $\lambda$ prediction (5). The predicted optimal $\lambda$ values are shared with the aggregator and then distributed to the source clients (Section 5.3.1).
Finally, using the optimal $\lambda$ values, the source clients collaboratively train final adaptive models in a federated manner. These final models are sent to the target client for inference (7).
@@ 2.3 Baselines
We compared the performance of freda against both wenda [6] and the non-adaptive model proposed by Horvath [20].
@@ 2.3.1 Wenda Baseline
Wenda has three external parameters: the weighting parameter $k$, the proportion of $L_{1}$ and $L_{2}$ penalties of the weighted elastic net $\alpha$, and the regularization parameter $\lambda$. The weighting parameter $k$ is inherent to the wenda method, and Handl et al. showed the performance of their method over different values of $k$ and select the best performing one. The parameters $\alpha$ and $\lambda$ are inherent to the standard elastic net and are typically optimized via cross-validation. Following the approach of Horvath, Hughey, and Butte [20, 21], Handl et al. treated $\alpha$ as a design choice and fixed it at $\alpha=0.8$. However, since the parameter $\lambda$ determines the strength of regularization, it cannot be globally set to a single value that performs well across different datasets [6]. Thus, Handl et al. proposed two variations of wenda, which differ in how the external parameter $\lambda$ is computed. One approach uses prior knowledge on tissue similarity to compute $\lambda$, and the other uses cross-validation on the training data to compute the optimal $\lambda$ value [6]. These variations are referred to as wenda-pn and wenda-cv, respectively.
Handl et al. [6] argue that using cross-validation to select the optimal $\lambda$ is counterintuitive in the context of unsupervised domain adaptation, as cross-validation cannot be performed on the target domain due to the absence of labels and must instead be conducted on the training (source) data.
Taking these considerations into account, we pick wenda-pn as our primary comparison baseline. In our experiments, we fix $\alpha=0.8$. For the weighting parameter $k$, we pick the best-performing value based on both our own experiments on wenda and the results reported by Handl et al. [6], which leads to $k=3$.
@@ 2.3.2 Non-Adaptive Baseline
Instead of using a simple elastic net, we adopt the non-adaptive baseline used in [6] and originally proposed by Horvath [20], which combines the elastic net with a least-squares fit. The idea is to first fit a standard elastic net and then apply a linear least-squares fit based only on features that obtained non-zero coefficients in the elastic net. This baseline was first proposed by Horvath [20] for age prediction from DNA methylation data, where he demonstrated that using an elastic net followed by a least-squares fit resulted in improved performance on his dataset. We refer to this non-adaptive method as en-ls.
2.4 Experimental Setup for Freda
We consider a distributed setting with source domain clients, a target client, and a helper party, referred to as the aggregator, which does not have any data. The labeled source domain data is distributed across multiple clients. We evaluate our framework in three different scenarios, with 2, 4, and 8 source domain clients, respectively.
@@ 2.4.1 Data Distribution
We assume that the source domain data is distributed uniformly at random among the source clients. Given that the DNA methylation dataset contains 1,866 samples in the training set, this results in approximately 933, 466, and 233 samples per source domain owner for the 2, 4, and 8 source client settings, respectively.
In order to showcase the robustness of our framework, we do not take the tissue types of the samples into consideration when distributing the source domain data among the source clients. Due to the unbalanced nature of DNA methylation data across tissues, as discussed in Section 2.2, the random distribution of samples results in some source clients to receive only a few samples from certain tissues, or in some cases, no samples from a particular tissue at all.
2.4.2 Federated Training Setup of the Weighted Elastic Net Models
The final weighted elastic net model is trained for 100 global iterations. During each iteration, the source clients update their local copy of the global model from the previous iteration for 20 epochs, after which the global model is updated with the help of the aggregator using secure aggregation.
During our experiments, we observed that the final model is highly sensitive to the learning rate. It fails to converge if the learning rate is too low and misses the global minimum if the learning rate is too high. To address this, we implement a learning rate decay based on the global iteration [22]. The source clients adjust the model’s learning rate using an exponential decay function, which gradually decreases the learning rate from an initial value to a final value over the course of the federated training process. The learning rate $\eta$ at global iteration $t$ is computed as follows:
@@ 2.4.2 Federated Training Setup of the Weighted Elastic Net Models
where $\eta(t)$ is the learning rate at iteration $t$, $T$ is the total number of iterations, and $t$ is the current iteration. For all of our experiments using freda, we set the initial learning rate $\eta_{0}$ to $1\times 10^{-4}$ and the final learning rate $\eta_{f}$ to $1\times 10^{-5}$.
2.4.3 External Parameters
Freda has the same three external parameters as wenda: the weighting parameter $k$, the proportion of $L_{1}$ and $L_{2}$ penalties of the weighted elastic net $\alpha$, and the regularization parameter $\lambda$. We consider $\alpha$ as a design choice and fix $\alpha=0.8$. For the regularization parameter $\lambda$, we adopt the prior knowledge approach (Section 5.3.1) proposed by Handl et al. [6].
For the tissue similarities required for the prior knowledge approach, we use the dataset from the GTEx consortium, which includes genotype and gene expression data across 42 human tissues [17]. Additionally, we follow the same steps described in [6] to compute tissue similarities. In the federated learning setting, only the owner of the target domain requires access to the tissue similarity information.
To evaluate the performance of our method, and following the evaluation strategy of Handl et al. for wenda-pn [6], we repeatedly split the test tissues into subsets: one for fitting the relationship between domain similarity and another for evaluation (Section 5.3.1). We iterate over all combinations of three tissues, each containing at least 20 samples for training, and then assess performance on the remaining tissues. Finally, we select the best-performing value for the weighting parameter $k$, which, based on our experiments is $k=3$.
@@ 2.5 Experiments
We compare the performance of freda on the DNA methylation dataset against wenda-pn and the non-private, non-adaptive baseline model en-ls, as described in Section 2.3. The main performance metric is the Mean Absolute Error (MAE) of the predicted chronological ages of the tissues. For wenda-pn, we calculate the MAE only on samples not used for fitting the tissue similarity-$\lambda$ relationship, reporting the mean and standard deviation across all splits. Similarly, for freda, we report the MAE exclusively for the target client’s tissues that were not part of the similarity-$\lambda$ fit, along with the mean and standard deviation over all splits.
@@
Figure 3: Predicted versus true chronological age for (a) en-ls and (b) wenda-pn with $k=3$. For en-ls, the predictions are averaged over 10 runs of 10-fold cross-validation, while for wenda-pn, the predictions are averaged over all splits of the test tissues where the tissue of interest was included in the evaluation set.
@@
Figures 3a and 3b display the predicted versus true ages for the samples of the target domain data, colored by tissue, for en-ls and wenda-pn, respectively. These plots clearly show that both methods perform well on most tissues, except for en-ls on cerebellum samples. Figure 3a reveals that the predicted ages for cerebellum samples are consistently lower than the true chronological ages. In contrast, Figure 3b demonstrates that the ages predicted by wenda-pn for cerebellum samples are much closer to the corresponding true ages. Furthermore, the predictions of wenda-pn on the remaining target domain tissues are quite similar to those of en-ls. This observation is confirmed by the quantitative comparison shown in Figure 2, where wenda-pn yields much lower errors than en-ls on cerebellum samples, while maintaining similar or better performance than en-ls on the remaining test tissues. Specifically, en-ls results in an MAE of $18.71\pm 7.13$ on cerebellum samples. In contrast, wenda-pn provides a substantial improvement in prediction performance on cerebellum samples, achieving an MAE of $7.63\pm 0.26$.
@@
Figure 4: Predicted versus true chronological age for freda with $k=3$ for (a) 2, (b) 4, and (c) 8 source parties. The predictions are averaged over all splits of the test tissues where the tissue of interest was included in the evaluation set, as well as over 5 different distributions for each setting.
@@
Focusing on the primary metric for evaluating the success of unsupervised domain adaptation, the performance of freda on cerebellum samples compared to our centralized baselines is shown in Figure 2. Despite running in a distributed setting, freda achieves comparable performance to wenda-pn on cerebellum samples for the 2 and 4 source domain scenarios, yielding an MAE of $7.99\pm 1.39$ and $8.64\pm 0.86$, respectively. In the 8 source domain scenario, freda yields an MAE of $10.77\pm 0.99$, which is only slightly worse than that of wenda-pn. However, even when the source domain data is distributed across 8 parties, freda still significantly outperforms the non-adaptive centralized baseline en-ls. This observation is further supported by the prediction results shown in Figure 4. Figure 4 shows the predicted versus true ages for the target domain data samples, colored by tissue, for freda under the 2, 4, and 8 source party settings, respectively. Comparing the predictive performance of freda with our baselines across Figures 3 and 4, it is clear that freda performs well across all tissues. Moreover, the consistently low predicted ages for cerebellum samples relative to their true chronological ages, as observed with en-ls (Fig. 3a), are not present in any of the freda settings. In comparison with wenda-pn (Fig. 3b), the results in Figure 4 show that freda performs similarly to wenda-pn, while effectively addressing the distribution shift associated with cerebellum samples, even in a distributed setting where the source domain data is spread across up to $8$ different parties.

### v2

# Privacy-Preserving Federated Unsupervised Domain Adaptation for Regression on Small-Scale and High-Dimensional Biological Data

@@ Abstract
Machine learning models often struggle with generalization in small, heterogeneous datasets due to domain shifts caused by variations in data collection and population differences. This challenge is particularly pronounced in biological data, where data is high-dimensional, small-scale, and decentralized across institutions. While federated domain adaptation methods (FDA) aim to address these challenges, most existing approaches rely on deep learning and focus on classification tasks, making them unsuitable for small-scale, high-dimensional applications. In this work, we propose freda, a privacy-preserving federated method for unsupervised domain adaptation in regression tasks. Unlike deep learning-based FDA approaches, freda is the first method to enable the federated training of Gaussian Processes to model complex feature relationships while ensuring complete data privacy through randomized encoding and secure aggregation. This allows for effective domain adaptation without direct access to raw data, making it well-suited for applications involving high-dimensional, heterogeneous datasets. We evaluate freda on the challenging task of age prediction from DNA methylation data, demonstrating that it achieves performance comparable to the centralized state-of-the-art method while preserving complete data privacy.
@@
Unsupervised Domain Adaptation, Privacy-Preserving Machine Learning, Small-Scale Biological Data
@@ 1 Introduction
Machine learning (ML) has rapidly become a powerful tool with applications across numerous fields, including computational biology and healthcare, where it has shown great potential in solving complex problems [1, 2, 3, 4]. However, collecting and labeling biological datasets is often challenging, costly, and time-consuming. As a result, many datasets in these fields are small-scale, unlabeled, and heterogeneous, often collected from different sources under varying environmental and experimental conditions—such as different laboratories, hospitals, or institutions [5]. These challenges introduce two critical issues: (1) data from different sources often exhibit distinct statistical distributions while lacking labeled samples, complicating direct model transfer; and (2) privacy regulations and the sensitive nature of biomedical data often restrict data sharing across institutions, necessitating collaborative learning approaches.
Unsupervised Federated Domain Adaptation (FDA) addresses these challenges by collaboratively aligning distributions between training and test data, referred to as source and target domains, respectively, without requiring direct data sharing [6]. The primary motivation for unsupervised FDA is the scarcity of labeled data in the target domain, making it impractical to train models from scratch. Most existing FDA methods aim to mitigate distributional differences between domains [7, 8, 9, 10]. While deep learning-based FDA methods have achieved success in computer vision [11, 12, 13, 14], their application to biological data remains limited due to high dimensionality and small sample sizes. Moreover, FDA research has predominantly focused on classification tasks, while regression-based approaches remain significantly underexplored despite their importance in biomedical applications [15, 16, 17].
In this context, we introduce freda (federated domain adaptation), a novel method for privacy-preserving, federated unsupervised domain adaptation in regression tasks. Unlike conventional deep learning-based approaches that struggle with data scarcity and high dimensionality, freda is the first method to leverage federated training of Gaussian Processes regressors (GPRs), enabling collaborating entities to model complex features without pooling their private data.
Gaussian Processes are particularly well-suited for feature modeling due to their probabilistic nature, providing not only point predictions but also uncertainty estimates in the form of Gaussian-distributed confidence intervals. This property is especially valuable in domain adaptation, where assessing prediction reliability is crucial when transferring knowledge across domains. However, like other kernel-based algorithms, GPRs require pairwise computation of data matrices, making it extremely challenging to train them when data is distributed across entities that cannot share raw samples.
To overcome this, freda introduces a novel combination of randomized encoding and secure aggregation, enabling federated training of Gaussian Processes while preserving complete data privacy. By facilitating robust feature modeling without direct access to raw data, freda is particularly well-suited for biological datasets, where privacy constraints, limited sample sizes, and data heterogeneity pose significant challenges.
We evaluate freda on a challenging benchmark task of age prediction from DNA methylation data. Our results demonstrate that freda achieves performance comparable to centralized methods while preserving complete data privacy. By providing a scalable, generalizable, and data-efficient solution for domain adaptation in biological datasets, freda enables secure collaboration across institutions. By addressing the challenges of small-scale, heterogeneous, and privacy-sensitive regression problems, our approach significantly expands the applicability of domain adaptation to real-world biomedical applications.
Our contributions are as follows:
  • We propose freda, the first method to enable privacy-preserving, federated unsupervised domain adaptation for regression tasks, specifically designed for small-scale, high-dimensional biological datasets.
  • Through a novel combination of randomized encoding and secure aggregation techniques, freda is the first method to enable the federated training of GPRs for effective feature modeling while ensuring complete data privacy.
  • We evaluate freda on the challenging task of age prediction from DNA methylation data, demonstrating that it effectively models complex feature relationships in small-scale, heterogeneous, and distributed biological datasets, achieving performance comparable to centralized approaches while preserving privacy.
@@ 2 Results
To evaluate the performance of our proposed method, we provide a benchmark on the problem of age prediction from DNA methylation data across multiple tissues. This section presents a detailed comparison of freda with existing baselines, highlighting its effectiveness in preserving data privacy while achieving competitive predictive accuracy.
@@ 2.1 Overview
We consider a scenario where the source domain data is distributed across multiple input parties, while the target domain data, held by one party, is unlabeled. The aim is to perform unsupervised domain adaptation on the target domain using labeled source data while preserving data privacy for all parties.
Figure 1: Overview of the freda framework, in which multiple source domain clients collaborate to perform domain adaptation on the target domain data with the assistance of an aggregator.
Figure 1 illustrates the four main phases of our framework:
  1. 1.
    Feature Model Training: Each feature in the dataset is modeled separately using federated hyper-parameter optimization and GPRs. In this phase, feature dependencies are learned in a federated and privacy-preserving manner using secure aggregation and randomized encoding.
  2. 2.
    Feature Weight Computation: The target client calculates feature confidence scores from the predicted distributions of the feature models and transforms them into feature weights. These weights are then shared with the aggregator and distributed to source clients.
  3. 3.
    Optimal Lambda Prediction: Source clients, with the help of the aggregator, federatively train multiple weighted elastic nets (WEN) with varying regularization parameters ($\lambda$). The optimal $\lambda$ values are identified by the target client and shared with all participants.
  4. 4.
    Final Adaptive Model Training: Using the optimal $\lambda$ values, source clients federatively train the final adaptive WEN models, which are then sent to the target client for inference.
This structured approach ensures both privacy preservation and robust performance across distributed data sources.
@@ 2.3 Baselines
We compared the performance of freda against the state-of-the-art unsupervised domain adaptation method for biological datasets proposed by Handl et al. [20] and a non-adaptive model proposed by Horvath [21].
@@ 2.3.1 Wenda Baseline
Wenda (weighted elastic net for domain adaptation) is a state-of-the-art method for unsupervised domain adaptation on small-scale high-dimensional biological datasets [20]. Like our approach, it leverages the dependency structure between inputs across source and target domains, penalizing discrepant features while emphasizing robust ones. Despite its strong performance over non-adaptive models, wenda assumes simultaneous access to both domains, hence it is only suitable for centralized settings.
Wenda has three key parameters: the weighting parameter $k$, the elastic net mixing parameter $\alpha$, and the regularization parameter $\lambda$. Following [20], we fix $\alpha=0.8$, while $\lambda$ is computed using prior knowledge on tissue similarity (wenda-pn), as cross-validation on the target domain is infeasible in an unsupervised setting. For $k$, we select $k=3$ based on both our experiments and prior work [20].
@@ 2.3.2 Non-Adaptive Baseline
For our non-adaptive baseline, we adopt the method proposed by Horvath [21], which combines the elastic net with a least-squares fit. The idea is to first fit a standard elastic net and then apply a linear least-squares fit based only on features that obtained non-zero coefficients in the elastic net. This baseline was first proposed by Horvath [21] for age prediction from DNA methylation data, where he demonstrated that using an elastic net followed by a least-squares fit resulted in improved performance on his dataset. We refer to this non-adaptive method as en-ls.
2.4 Setup for Freda
We consider a distributed setting with multiple source domain clients, a target client, and an aggregator, which has no data. The labeled source domain data is distributed across 2, 4, or 8 source clients, and we evaluate freda in each of these settings.
@@ 2.4.1 Data Distribution
Source domain data is assigned uniformly at random among the source clients. Given that the DNA methylation dataset contains 1,866 training samples, each client receives approximately 933, 466, or 233 samples in the 2, 4, and 8-client settings, respectively.
To assess the robustness of freda, we do not consider tissue types when distributing data. Due to the inherent imbalance of DNA methylation data across tissues (Section 2.2), this results in some clients having only a few or no samples from certain tissues.
2.4.2 Setup for Weighted Elastic Net Models
The final weighted elastic net model is trained for 100 global iterations. In each iteration, source clients update their local model for 20 epochs before the aggregator securely updates the global model.
Furthermore, we apply an exponential learning rate decay based on global iterations [22]:
@@ 2.4.2 Setup for Weighted Elastic Net Models
where $\eta(t)$ is the learning rate at iteration $t$, $T$ is the total number of iterations, and $t$ is the current iteration. We set the initial and final learning rates to $\eta_{0}=1\times 10^{-4}$ and $\eta_{f}=1\times 10^{-5}$, respectively.
2.4.3 Parameter Selection in Freda
Freda has three key parameters: the weighting parameter $k$, the elastic net mixing parameter $\alpha$, and the regularization parameter $\lambda$. We fix $\alpha=0.8$ as a design choice and determine $\lambda$ using the prior knowledge approach (Section 5.3.1) proposed by Handl et al. [20].
For tissue similarity calculations, we use data from the GTEx consortium, which provides genotype and gene expression data across 42 human tissues [23], following the methodology in [20]. In the federated setting, only the target domain owner requires access to tissue similarity information.
To evaluate performance, we follow the evaluation strategy of Handl et al. for wenda-pn [20], iteratively splitting test tissues into subsets: one for fitting domain similarity relationships and another for evaluation (Section 5.3.1). We iterate over all three-tissue combinations with at least 20 training samples, assessing performance on the remaining tissues. Based on our experiments, the optimal weighting parameter is $k=3$.
@@ 2.5 Experiments
We compare the performance of freda against wenda-pn and the non-private, non-adaptive baseline model en-ls, as described in Section 2.3. The main performance metric is the Mean Absolute Error (MAE) of the predicted chronological ages of the tissues. For wenda-pn, we calculate the MAE only on samples not used for fitting the tissue similarity-$\lambda$ relationship, reporting the mean and standard deviation across all splits. Similarly, for freda, we report the MAE exclusively for the target client’s tissues that were not part of the similarity-$\lambda$ fit, along with the mean and standard deviation over all splits.
@@
Figure 3: Predicted versus true chronological age under various settings. Figures (a), (b), and (c) correspond to freda with $k=3$ for 2, 4, and 8 source parties, respectively. Predictions are averaged over all splits where the tissue of interest was included in the evaluation set, as well as over 5 different distributions for each setting. Panels (d) and (e) correspond to en-ls and wenda-pn, respectively. For en-ls, predictions are averaged over 10 runs of 10-fold cross-validation, while for wenda-pn, predictions are averaged over all splits where the tissue of interest was included in the evaluation set.
Figure 3 show the predicted versus true ages for the samples of the target domain data, colored by tissue, for freda with $k=3$ for 2, 4, and 8 source parties, en-ls and wenda-pn, respectively. From Figures 3d and 3e, we can clearly see that both centralized methods perform well on most tissues, except for en-ls on cerebellum samples. As shown in Figure 3d, the ages predicted by en-ls for cerebellum samples are consistently lower than the true chronological ages. In contrast, Figure 3e demonstrates that wenda-pn achieves much closer alignment between the predicted and true ages for cerebellum samples.
Additionally, for the remaining target domain tissues, the predictions of wenda-pn are comparable to those of en-ls, as confirmed by the quantitative results in Figure 2. Wenda-pn not only yields significantly lower errors than en-ls on cerebellum samples but also maintains similar or better performance on other test tissues. Specifically, on cerebellum samples, en-ls produces a mean absolute error (MAE) of $7.63\pm 0.26$. These results highlight the significance of improving prediction performance on cerebellum samples without a drop in performance on other tissues.
@@
Focusing on the primary metric for evaluating the success of unsupervised domain adaptation, the performance of freda on cerebellum samples compared to our centralized baselines is shown in Figure 2. Despite running in a distributed setting, freda achieves comparable performance to wenda-pn on cerebellum samples for the 2 and 4 source domain scenarios, yielding an MAE of $7.99\pm 1.39$ and $8.64\pm 0.86$, respectively. In the 8 source domain scenario, freda yields an MAE of $10.77\pm 0.99$, which is only slightly worse than that of wenda-pn. However, even when the source domain data is distributed across 8 parties, freda still significantly outperforms the non-adaptive centralized baseline en-ls. This observation is further supported by the prediction results shown in Figure 3. Figures 3a, 3b, and 3c shows the predicted versus true ages for the target domain data samples, colored by tissue, for freda under the 2, 4, and 8 source party settings, respectively. Comparing the predictive performance of freda with our baselines, it is clear that freda performs well across all tissues. Moreover, the consistently low predicted ages for cerebellum samples relative to their true chronological ages, as observed with en-ls, are not present in any of the freda settings. In comparison with wenda-pn, the results in Figure 3 show that freda performs similarly to wenda-pn, while effectively addressing the distribution shift associated with cerebellum samples, even in a distributed setting where the source domain data is spread across up to $8$ different parties.